# Recent Advances in Speech Language Models: A Survey

### v3

@@ I Introduction
Figure 1: Architectures of the “ASR + LLM + TTS” framework and a SpeechLM. We emphasize that, for SpeechLM, the same content can be used across both speech and text modalities, meaning that any input modality will yield any output modality of the same results. This intentional repetition of input/output contents in the figure highlights this point.
@@
$ms$ Mel-spectrogram of a speech audio waveform
@@
Contributions. In this survey, we present the first comprehensive overview of recent endeavors in constructing SpeechLMs. We explore the various components that constitute their architecture (section III) and the training recipes (table III) involved in their development. we aim to elucidate the current state of the field by analyzing these models from the above perspectives. Additionally, we survey the downstream applications of SpeechLMs (section V), classify metrics to evaluate SpeechLMs (section VI), discuss the challenges encountered in this rapidly evolving area, and outline promising future research directions that could drive further advancements in SpeechLM technology (section VII). Our contributions are summarized as follows:
@@ III-A Speech Tokenizer
GSLM [37], TWIST [38], SpeechGPT [8], AudioPaLM [39], OmniFlatten [40], SLAM-Omni [41] , color=lightgreen, fill=lightgreen!15, text width=5.3cm, text=black], [ Paralinguistic Token:
@@
(section IV-C), color=lightgreen, fill=lightgreen!15, text width=2.1cm, text=black [ dGSLM [4], Parrot [45], LSLM [51], VITA [55], Moshi [9], Mini-Omni 2 [52], OmniFlatten [40], SALMONN-omni [56] , color=lightgreen, fill=lightgreen!15, text width=5.3cm, text=black], ] [Silence Mode
(section IV-C), color=lightgreen, fill=lightgreen!15, text width=2.1cm, text=black [ VITA [55] , color=lightgreen, fill=lightgreen!15, text width=5.3cm, text=black], ] ] ] ]
@@
Speech tokenizer is the first component in SpeechLMs, which encodes continuous audio signals (waveforms) into tokens. Speech tokenizer aims to capture essential features of the audio while reducing its dimensionality and allows the audio input to be effectively processed by a language model for autoregressive generation. Speech tokenizer operates by encoding the audio segment by segment, producing two possible types of tokens (features): discrete tokens and continuous tokens. Discrete tokens (Section IV-A1) use a specific index to represent each speech segment, while continuous tokens (Section IV-A2) use an embedding to represent the segment [^2] For simplicity, we unify both types of representations as ”tokens”. Both token types can be utilized as input for a language model in autoregressive modeling. In this section, we categorize speech tokenizers based on their focus on modeling different aspects of the raw audio.
@@ III-A1 Semantic Understanding Objective
[…]
The key design choices lie in how to effectively encode (and quantize) speech into tokens. Wav2vec 2.0 [17] uses a convolutional encoder followed by a product quantization module [57] to discretize the continuous waveform. Then, a portion of the quantized representations is masked and modeled using a contrastive loss. W2v-BERT [19] is built upon wav2vec 2.0 and proposes to use Masked Language Modeling (MLM) loss [58] in addition to contrastive loss. Similarly, HuBERT [20] uses the k-means algorithm to cluster the speech utterances into a number of hidden units, and then perform MLM to predict the target hidden units from the masked speech utterances. To better align the representation of text and speech modalities, Google USM [23] utilizes text-injection loss [59] at the second pre-training stage to improve the performance and robustness of the downstream tasks. WavLM [24] adds the speech denoising objective during pre-training. While the majority of speech tokenizer studies focus on semantic-related tasks such as ASR and TTS, WavLM shows that speech denoising can boost the performance of non-semantic tasks such as speaker verification and speech separation. Section V presents a full list of downstream tasks.
@@ III-A3 Mixed Objective
TABLE II: Summarization of the architectural choice of speech tokenizer, language model, and vocoder in popular SpeechLMs. “-” represents non-existence or not indicated, * means the architecture is mainly based on the written one, “A, B” means the authors experimented with both A and B as the component, and “A + B” means “A” and “B” are combined to serve as the component.
@@
SLAM-Omni [41] Whisper Encoder  [12] + Linear Projector Qwen2 [28] -
OmniFlatten [40] CosyVoice Encoder [61] Qwen2 CosyVoice Decoder
@@
Zeng et al. [70] Whisper + ASR GLM [29] CosyVoice
Parrot [45] VQ-VAE LLaMA-3, Mistral, Gemma 2 HiFi-GAN
@@ IV-A1 Discrete Features
Another type is acoustic tokens, which are tokens aiming to capture essential acoustic features to reconstruct high-fidelity speech, primarily obtained from neural audio codec models (Section III-A2). Some studies directly model the codec tokens in a language model, which is often regarded as Codec Language Models (CodecLMs). For example, Viola [43] trains a CodecLM capable of performing ASR, TTS, and Machine Translation. Parrot [45] trains on VQ-VAE [60] tokens for modeling dual-channel spoken dialogue data.
@@ IV-B3 Language Model Post-Alignment
IV-C Speech Generation Paradigm
Most approaches covered in earlier sections follow the traditional generation paradigm of SpeechLMs, which involves taking a predefined input sequence and generating a complete response. However, this approach does not reflect the natural flow of voice interactions. For instance, during a conversation, one person may interrupt another, switching from listening to speaking. Additionally, a person might choose not to respond if the other is engaged in a conversation with someone else. Based on these observations, we identify two key aspects of advanced voice interaction skills for SpeechLMs: real-time interaction and interactive period recognition.
Real-time Interaction refers to the capability of SpeechLMs to engage with users instantaneously. This interaction consists of two key components. 1) User Interruption: SpeechLMs should be able to be interrupted by users and should respond appropriately to new instructions provided during the conversation. 2) Simultaneous Response: SpeechLMs should be capable of generating responses while the user is still speaking. This requires the model to effectively perform speech understanding (processing input) and speech generation (producing output) simultaneously. Some literature refers to the ability as full-duplex modeling. dGSLM [4] introduces a dual-transformer architecture to model two-speaker dialogues, using one transformer to handle speech from each speaker. A cross-attention transformer layer is included to capture the interactions between the speakers’ content. However, most approaches rely on a single language model architecture to manage both channels in a dialogue. Parrot [45] introduces the ”next-token-pair prediction” approach, which uses a single decoder-only transformer to model both channels simultaneously by predicting tokens for both channels at each time step. Moshi [9] concatenates the user input channel and the model response channel and employs an RQ-Transformer that first processes data along the temporal dimension and then along the channel dimension. LSLM [51] utilizes a decoder-only Transformer focusing on modeling one speaker’s speech in the dialogue. This model incorporates a streaming SSL encoder that continuously processes input from the listening channel and fuses its embeddings with those from the speaking channel.
@@ V Downstream Applications
Speaker Diarization [An audio file of the meeting] Please provide the speaker diarization output in the following format: [Start Time] - [End Time]: Speaker [Speaker]. Sure. Here is the speaker diarization result: $0$ 00:15: Alice, 00:16 - 00:40: Bob, 00:41 - 01:00: Carol, 01:01 - 01:20: Alice, 01:21 - 01:45: Bob and Carol (overlapping speech), 01:46 - 02:00: Carol.
@@ VI Evaluations
TABLE VI: A summary of popular benchmarks for the evaluation of SpeechLMs. I/O, $A$, and $T$ represent input/output format, audio, and text, respectively.
@@
VoiceBench [133] Downstream 8 Speech $A\rightarrow A$
@@ VI-A Automatic (Objective) Evaluation
Paralinguistic Evaluation. In contrast to linguistic evaluation, paralinguistic evaluation focuses on the non-verbal aspects of communication that accompany speech. Some works choose to utilize paralinguistic tokens alongside semantic tokens to enhance the paralinguistic abilities of SpeechLMs [42, 5], so one way is to evaluate the paralinguistic tokens. pGSLM [42] measures the correctness, consistency, and expressiveness of the prosodic tokens. Correctness evaluates the model’s ability to generate accurate prosodic profiles by calculating the minimal mean absolute error (min-MAE) of the prosodic tokens from 20 generated samples against the prosodic tokens from the reference, consistency is assessed through the Pearson correlation between the mean values of the prompt prosodic and its generated continuation prosodic tokens, and expressiveness is measured by the standard deviation of the generated prosody token values, with the expectation that it matches the variability of the ground truth. We note that the same metrics can also be applied to other paralinguistic tokens. Instead of evaluating from the token level, SPIRIT-LM [5] proposes to measure on the perceptual level. They introduced a speech-text sentiment preservation benchmark (STSP), which requires the model to generate a text or speech sequence of tokens that preserves the sentiment of the prompt. A sentiment classifier is used to assess the sentiment in the generated speech. It should be noted that although they only apply the preservation approach on sentiment, this idea can be generalized to other paralinguistic features, such as timbre or prosody.
@@
Generation Quality and Diversity. Quality and diversity are two crucial aspects of model generation. Typically, there is a trade-off between these dimensions when sampling model responses at different temperatures, so GSLM [37] suggests using the Area Under the Curve (AUC) with various temperature values. Specifically, AUC on perplexity and VERT are employed to assess these factors, where VERT represents the geometric mean of the ratio of k-grams in the generated speech that appears at least once. Additionally, the ChatGPT score can be utilized to evaluate the quality of the generated speech. In this process, the generated speech is transcribed using state-of-the-art ASR models and then sent to ChatGPT for quality (and diversity) assessment.
@@
Real-Time Interaction Evaluation. Real-time interaction evaluation involves assessing the ability of SpeechLMs to interact in real-time, which is crucial for those models that facilitate full-duplex communication. Current research focuses on evaluating the naturalness and usefulness of speech generated in real-time. dGSLM [4] examines the naturalness of dialogues between two speakers by introducing different turn-taking events, such as speech segments (Inter-Pausal Unit, IPU), pauses within speech (pause), pauses between speeches (gaps), and overlapping speech (overlap). The generated speech is considered more natural if the statistics of these turn-taking events closely resemble those found in human dialogues. Another method involves speech continuation, where the generated speech is considered more natural if the turn-taking statistics of the speech prompts align closely with those of the subsequent continuations. Additionally, evaluating SpeechLM’s usefulness as an AI assistant in real-time interaction is crucial. The Parrot model [45] suggests using reflective pauses and interruptions to assess usefulness, where reflective pause evaluates the SpeechLM’s ability to remain silent while the user speaks, and interruption measures the SpeechLM’s capability to cease speaking when interrupted.
@@
Downstream Evaluation. Downstream evaluation refers to evaluating the ability of SpeechLMs to perform specific tasks, such as ASR, TTS, Speaker Identification, etc. The evaluation can be performed on pre-trained models by adding few-shot example(s) at the start of the prompt or on the instruction-tuned models by directly instructing them to do so. Several benchmarks compile a broad array of downstream tasks to provide a comprehensive assessment of SpeechLMs. SUPERB [123] includes a variety of speech understanding tasks. SD-Eval [130] assesses the paralinguistic understanding of SpeechLMs using the emotion, age, environment, and age classification tasks. SALMON [132] tests the SpeechLMs’ ability to generate speech with consistent paralinguistic and environmental characteristics. Dynamic-SUPERB [131], MMAU [127], AirBench [129], and AudioBench [128] extend beyond traditional speech tasks to include sound and/or music-related challenges as well.
@@
While these benchmarks offer comprehensive coverage of audio-related tasks, they primarily require models to respond in text, which creates a barrier to end-to-end speech interaction evaluation. To address this limitation, newer benchmarks have emerged that support speech output. VoxEval [134] focuses on benchmarking the knowledge understanding capabilities of SpeechLMs, providing question-answer pairs in comprehensive subjects in speech format, along with an evaluation pipeline tailored for speech output. Additionally, it presents questions under various input audio conditions for evaluating the model’s robustness. Similarly, VoiceBench [133] supports speech output while focusing on assessing general knowledge, instruction following, and response safety of SpeechLMs.

### v4

@@ I Introduction
Figure 1: Architectures of the “ASR + LLM + TTS” framework and a SpeechLM. We emphasize that, for SpeechLM, the same content can be used across both speech and text modalities, meaning that any input modality can yield any output modality of the same results. This intentional repetition of input/output contents in the figure highlights this point.
@@
msms Mel-spectrogram of a speech audio waveform
@@
Contributions. In this survey, we present the first comprehensive overview of recent endeavors in constructing SpeechLMs. We explore the various components that constitute their architecture (Section III) and the training recipes (Section III) involved in their development. we aim to elucidate the current state of the field by analyzing these models from the above perspectives. Additionally, we survey the downstream applications of SpeechLMs (Section V), classify metrics to evaluate SpeechLMs (Section VI), discuss the challenges encountered in this rapidly evolving area, and outline promising future research directions that could drive further advancements in SpeechLM technology (Section VII). Our contributions are summarized as follows:
@@
Connections with other surveys. Several surveys have concentrated on traditional speech and audio technologies, such as Spoken Language Understanding (SLU) [16], audio and speech Self-Supervised Learning (SSL) [17, 18], and the integration of speech with other modalities [19, 20]. With the rapid advancements in LLMs, some studies review single-modal [21, 22] and multi-modal [23, 24, 25] LLMs. Additionally, there are surveys that explore the overlap between the audio modality and LLMs. Latif et al. [26] examine LLMs in audio processing, Peng et al. [27] review SpeechLLMs within the SLU domain, and Ji et al. [28] focus on spoken dialogue systems that encompass speech, sound, and music.
@@ III-A Speech Tokenizer
Figure 3: Illustration of the three types of speech tokenizers.
@@
GSLM [50], TWIST [51], SpeechGPT [8], AudioPaLM [52], OmniFlatten [53], SLAM-Omni [54], GLM-4-Voice [55] , color=lightgreen, fill=lightgreen!15, text width=5.3cm, text=black], [ Paralinguistic Token:
@@
(section IV-C), color=lightgreen, fill=lightgreen!15, text width=2.1cm, text=black [ dGSLM [4], NTPP [69], LSLM [65], VITA [70], Moshi [9], Mini-Omni 2 [66], OmniFlatten [53], SALMONN-omni [71] , color=lightgreen, fill=lightgreen!15, text width=5.3cm, text=black], ] [Interactive Period Recognition
(section IV-C), color=lightgreen, fill=lightgreen!15, text width=2.1cm, text=black [ VITA [70], MiniCPM-o 2.6 [72], FlexDuo [73] , color=lightgreen, fill=lightgreen!15, text width=5.3cm, text=black], ] ] ] ]
@@
Speech tokenizer is the first component in SpeechLMs, which encodes continuous audio signals (waveforms) into tokens. Speech tokenizer aims to capture essential features of the audio while reducing its dimensionality and allows the audio input to be effectively processed by a language model for autoregressive generation. Speech tokenizer operates by encoding the audio segment by segment, producing two possible types of tokens (features): discrete tokens and continuous tokens. Discrete tokens (Section IV-A1) use a specific index to represent each speech segment, while continuous tokens (Section IV-A2) use an embedding to represent the segment [^2] For simplicity, we unify both types of representations as ”tokens”. Both token types can be utilized as input for a language model in autoregressive modeling. In this section, we divide speech tokenizers into three categories based on their focus on modeling different aspects of the raw audio. Figure 3 illustrates the three types of speech tokenizers.
@@ III-A1 Semantic Understanding Objective
A semantic understanding speech tokenizer typically comprises a speech encoder and a quantizer. The speech encoder encodes the essential information from the waveform into continuous embeddings. Then, a quantizer is typically incorporated to convert continuous embeddings into discrete indexes. Let $f_{E}(\cdot)$ denote the speech encoder parameterized by $\theta_{f_{E}}$, we have $\textbf{v}=f_{E}(\textbf{a};\theta_{f_{E}})$, where $\textbf{v}=(v_{1},v_{2},\ldots,v_{P})$ represents the encoded embeddings. Since v is still continuous, a quantizer $d(\cdot)$ is utilized to discretize the embeddings. Depending on different design choices, the speech tokens $\textbf{s}=(s_{1},s_{2},\ldots,s_{P})$ can either be derived from a or v. Therefore, we have $\textbf{s}=d(\textbf{v};\theta_{d})$ or $\textbf{s}=d(\textbf{a};\theta_{d})$ for discrete tokens and $\textbf{s}=\textbf{v}$ for continuous tokens. After that, s can be used to train the speech tokenizer as a target label (such as masking $\textbf{a}_{\text{mask}}\subset\textbf{a}$ and reconstructing its corresponding label $\textbf{s}_{\text{mask}}\subset\textbf{s}$ [33]) or to train the following language model.
The key design choices lie in how to effectively encode and/or quantize speech into tokens. Wav2vec 2.0 [30] uses a convolutional encoder followed by a product quantization module [74] to discretize the continuous waveform. Then, a portion of the quantized representations is masked and modeled using a contrastive loss. W2v-BERT [32] is built upon wav2vec 2.0 and proposes to use Masked Language Modeling (MLM) loss [75] in addition to contrastive loss. Similarly, HuBERT [33] uses the k-means algorithm to cluster the speech utterances into a number of hidden units, and then perform MLM to predict the target hidden units from the masked speech utterances. To better align the representation of text and speech modalities, Google USM [36] utilizes text-injection loss [76] at the second pre-training stage to improve the performance and robustness of the downstream tasks. WavLM [37] adds the speech denoising objective during pre-training. While the majority of speech tokenizer studies focus on semantic-related tasks such as ASR and TTS, WavLM shows that speech denoising can boost the performance of non-semantic tasks such as speaker verification and speech separation. Section V presents a full list of downstream tasks.
@@ III-A3 Mixed Objective
In the following paragraphs, we present the notations for three representative speech tokenizers, each exemplifying a distinct category. We examine HuBERT as a semantic objective tokenizer, Encodec as an acoustic objective tokenizer, and SpeechTokenizer as a mixed objective tokenizer.
HuBERT. As a representative semantic objective tokenizer, HuBERT [33] employs a feature encoder $f_{E}$ to transform raw audio waveforms $\mathbf{a}$ into continuous embeddings $\mathbf{v}$, i.e., $f_{E}(\mathbf{a};\theta_{f_{E}})=\mathbf{v}$. These embeddings are then quantized into discrete speech tokens $\mathbf{s}$ via k-means clustering of MFCC features, denoted as $d(\mathrm{MFCC}(\mathbf{a});\theta_{d})=\mathbf{s}$. The model is trained with a masked prediction objective, which seeks to maximize the likelihood of the correct token at masked positions:
$\mathcal{L}(\theta)=\mathbb{E}_{\mathbf{a}\sim\mathcal{D}}\left[\sum_{i\in\mathcal{M}}-\log p(s_{i}\mid\mathbf{v}_{\setminus\mathcal{M}};\theta)\right],$ (2)
where $\mathcal{M}$ denotes the masked indices. HuBERT further refines its speech tokens iteratively, updating the encoder and discretizer parameters at each step as
$\mathbf{s}^{(n+1)}=d(f_{E}(\mathbf{a};\theta_{f_{E}}^{(n)});\theta_{d}^{(n)}).$ (3)
This iterative process enables the learning of increasingly meaningful speech representations.
Encodec. As a representative acoustic objective tokenizer, EnCodec [49] employs a convolutional encoder-decoder architecture with residual vector quantization (RVQ). The encoder $f_{E}$ maps the raw audio waveform $\mathbf{a}$ to continuous embeddings $\mathbf{v}$, i.e., $\mathbf{v}=f_{E}(\mathbf{a};\theta_{f_{E}})$. These embeddings are then discretized using a multi-stage RVQ, where each stage $r$ quantizes the residual from the previous stage:
$\mathbf{s}=d(\mathbf{v};\theta_{d})=\big{(}$ $d_{1}(\mathbf{v};\theta_{d_{1}}),\ d_{2}(\mathbf{v}-\hat{\mathbf{v}}_{1};\theta_{d_{2}}),$
$\ldots,\ d_{R}(\mathbf{v}-\sum_{r=1}^{R-1}\hat{\mathbf{v}}_{r};\theta_{d_{R}})\big{)},$ (4)
with $\hat{\mathbf{v}}_{r}$ denoting the quantized embedding at stage $r$. The decoder $f_{D}$ reconstructs the audio waveform from the quantized tokens, $\hat{\mathbf{a}}=f_{D}(\mathbf{s};\theta_{f_{D}})$. This design enables EnCodec to produce discrete acoustic tokens that retain high-fidelity audio information suitable for downstream modeling.
SpeechTokenizer. As a representative mixed objective tokenizer, SpeechTokenizer [34] combines semantic and acoustic objectives by leveraging both HuBERT and residual vector quantization (RVQ) mechanisms. The encoder $f_{E}$ first transforms the input audio waveform $\mathbf{a}$ into continuous embeddings $\mathbf{v}$, i.e., $\mathbf{v}=f_{E}(\mathbf{a};\theta_{f_{E}})$. Discretization is performed via a multi-stage RVQ. The discretization process uses a multi-stage RVQ, which operates similarly to Encodec, except that the first RVQ stage distills tokens derived from HuBERT, while the subsequent stages quantize the residuals. This hybrid approach enables SpeechTokenizer to capture both high-level semantic and low-level acoustic information for robust speech representation learning.
TABLE II: Summarization of the architectural choice of speech tokenizer, language model, and vocoder in popular SpeechLMs. “-” represents non-existence or not indicated, * means the architecture is mainly based on the written one, “A, B” means the authors experimented with both “A” and “B” as the component, and “A + B” means “A” and “B” are combined to serve as the component.
@@
Kimi-Audio [78] Whisper Encoder [12] + Linear Projector Qwen2.5 [79] BigVGAN [80]
Qwen2.5-Omni [81] Whisper Qwen2.5 Talker + Codec Decoder [81]
Minmo [82] SenseVoice [83] Qwen2.5 CosyVoice 2 [84]
Lyra [85] Whisper [12] Qwen2-VL [86] HuBERT + HiFi-GAN
Flow-Omni [87] Whisper Encoder + Linear Projector Qwen2 [41] Flow Matching (Transformer + MLP) + HiFi-GAN
SLAM-Omni [54] Whisper Encoder + Linear Projector Qwen2 [41] -
OmniFlatten [53] CosyVoice Encoder [88] Qwen2 CosyVoice Decoder [88]
@@
Zeng et al. [97] Whisper + VQ GLM [42] CosyVoice
NTPP [69] VQ-VAE LLaMA-3, Mistral, Gemma 2 HiFi-GAN
@@
Park et al. [104] AV-HuBERT [105] OPT HiFi-GAN
USDM [106] XLS-R [107] Mistral Voicebox [108]
@@
ICoT [icot] SpeechTokenizer LLaMA-2 SoundStorm
AnyGPT [111] SpeechTokenizer LLaMA-2 SoundStorm
@@ III-C3 Other Types of Vocoder
In the following paragraph, we present the notation of HiFi-GAN [47] as it is the most used vocoder in SpeechLMs. HiFi-GAN synthesizes high-fidelity audio waveforms from mel-spectrograms or speech tokens using a generator-discriminator framework. The generator $G(\mathbf{s};\theta_{G})$ maps a sequence of speech tokens $\mathbf{s}$ to an output audio waveform $\mathbf{a}$, i.e.,
$\mathbf{a}=Vo(\mathbf{s};\theta_{Vo})=G(\mathbf{s};\theta_{G}),$ (13)
where $Vo$ denotes the vocoder function and $\theta_{Vo}=\theta_{G}$ are its parameters. HiFi-GAN employs multi-period and multi-scale discriminators, $D_{MPD}(\mathbf{a};\theta_{MPD})$ and $D_{MSD}(\mathbf{a};\theta_{MSD})$, to distinguish real from generated audio during adversarial training. At inference, only the generator $G$ is used to efficiently reconstruct speech waveforms.
@@ IV-A1 Discrete Features
Another type is acoustic tokens, which are tokens aiming to capture essential acoustic features to reconstruct high-fidelity speech, primarily obtained from neural audio codec models (Section III-A2). Some studies directly model the codec tokens in a language model, which is often regarded as Codec Language Models (CodecLMs). For example, Viola [57] trains a CodecLM capable of performing ASR, TTS, and Machine Translation. NTPP [69] trains on VQ-VAE [77] tokens for modeling dual-channel spoken dialogue data.
@@ IV-B3 Language Model Post-Alignment
IV-C Speech Interaction Paradigm
Most approaches covered in earlier sections follow the traditional generation paradigm of SpeechLMs, which involves taking a predefined input sequence and generating a complete response. However, this approach does not reflect the natural flow of voice interactions. For instance, during a conversation, one person may interrupt another, switching from listening to speaking. Additionally, a person might choose not to respond if the other is engaged in a conversation with someone else. Based on these observations, we identify two key aspects of advanced speech interaction skills for SpeechLMs: real-time interaction and interactive period recognition.
Real-time Interaction of SpeechLMs involves the advanced handling of conversation data from two or more people, and it can be understood through several progressive stages. The initial stage is the adoption of streaming tokenizers and vocoders, which eliminate the need for the language model to wait for complete speech encoding before processing. This architecture enables immediate, low-latency responses to user queries, marking a significant improvement over the traditional interaction paradigm. Nonetheless, while this streaming approach supports basic real-time interaction, it remains insufficient for capturing the more sophisticated interaction patterns observed in natural conversation. The next frontier is full-duplex modeling, which allows SpeechLMs to support simultaneous bidirectional communication—specifically, the ability to handle interruptions initiated by either the user or the model. It mainly includes two features: 1) User interruption, where models can be interrupted and respond appropriately to new instructions during a conversation, and 2) Simultaneous response, enabling models to process input and generate output concurrently. Achieving this requires the joint modeling of both user and model audio streams. dGSLM [4] employs a separate transformer for each participant in two-speaker dialogues, with cross-attention layers capturing speaker interactions. Most methods, however, rely on a single language model. NTPP [69] employs a “next-token-pair prediction” approach with a decoder-only Transformer to predict tokens for both channels. Moshi [9] concatenates user input and model response channels data, using an RQ-Transformer to process the data together. LSLM [65] focuses on modeling one speaker’s speech using a decoder-only Transformer, integrating a streaming SSL encoder to fuse listening and speaking channel embeddings.
@@ V Downstream Applications
Speaker Diarization [An audio file of the meeting] Please provide the speaker diarization output in the following format: [Start Time] - [End Time]: Speaker [Speaker]. Sure. Here is the speaker diarization result: 00:00 - 00:15: Alice, 00:16 - 00:40: Bob, 00:41 - 01:00: Carol, 01:01 - 01:20: Alice, 01:21 - 01:45: Bob and Carol (overlapping speech), 01:46 - 02:00: Carol.
@@ VI Evaluations
TABLE VI: A summary of popular benchmarks for the evaluation of SpeechLMs. I/O, AA, and $T$ represent input/output modality, audio, and text, respectively.
@@
VoxDialogue [164] Downstream 12 Speech, Sound, Music $A\rightarrow T$
@@
VoiceBench [166] Downstream 8 Speech $A\rightarrow T$
@@ VI-A Automatic (Objective) Evaluation
Paralinguistic Evaluation. In contrast to linguistic evaluation, paralinguistic evaluation focuses on the non-verbal aspects of communication that accompany speech. Some works choose to utilize paralinguistic tokens alongside semantic tokens to enhance the paralinguistic abilities of SpeechLMs [56, 5], so one way is to evaluate the paralinguistic tokens. pGSLM [56] measures the correctness, consistency, and expressiveness of the prosodic tokens. Correctness evaluates the model’s ability to generate accurate prosodic profiles by calculating the minimal mean absolute error (min-MAE) of the prosodic tokens from 20 generated samples against the prosodic tokens from the reference, consistency is assessed through the Pearson correlation between the mean values of the prompt prosodic and its generated continuation prosodic tokens, and expressiveness is measured by the standard deviation of the generated prosody token values, with the expectation that it matches the variability of the ground truth. We note that the same metrics can also be applied to other paralinguistic tokens. Instead of evaluating from the token level, SPIRIT-LM [5] proposes to measure on the perceptual level. They introduced a speech-text sentiment preservation benchmark (STSP), which requires the model to generate a text or speech sequence of tokens that preserves the sentiment of the prompt. A sentiment classifier is used to assess the sentiment in the generated speech. It should be noted that although they only apply the preservation approach to sentiment, this idea can be generalized to other paralinguistic features, such as timbre or prosody.
@@
Generation Quality and Diversity. Quality and diversity are two crucial aspects of model generation. Typically, there is a trade-off between these dimensions when sampling model responses at different temperatures, so GSLM [50] suggests using the Area Under the Curve (AUC) with various temperature values. Specifically, AUC on perplexity and VERT are employed to assess these factors, where VERT represents the geometric mean of the ratio of k-grams in the generated speech that repeat at least once. Additionally, the ChatGPT score can be utilized to evaluate the quality of the generated speech. In this process, the generated speech is transcribed using state-of-the-art ASR models and then sent to ChatGPT for quality (and diversity) assessment.
@@
Real-Time Interaction Evaluation. Real-time interaction evaluation involves assessing the ability of SpeechLMs to interact in real-time, which is crucial for those models that facilitate streaming or full-duplex interaction. Current research focuses on evaluating the naturalness and usefulness of speech generated in real-time. dGSLM [4] examines the naturalness of dialogues between two speakers by introducing different turn-taking events, such as speech segments (Inter-Pausal Unit, IPU), pauses within speech (pause), pauses between speeches (gaps), and overlapping speech (overlap). The generated speech is considered more natural if the statistics of these turn-taking events closely resemble those found in human dialogues. Another method involves speech continuation, where the generated speech is considered more natural if the turn-taking statistics of the speech prompts align closely with those of the subsequent continuations. Recently, Full Duplex Bench [168] serves as a benchmark to evaluate various turn-taking abilities in full duplex SpeechLMs, and Talking Turns [169] evaluate turn-taking events by training a neural network model to predict the events outputted by the full duplex SpeechLMs. Additionally, evaluating SpeechLM’s usefulness as an AI assistant in real-time interaction is crucial. The NTPP model [69] suggests using reflective pauses and interruptions to assess usefulness, where reflective pause evaluates the SpeechLM’s ability to remain silent while the user speaks, and interruption measures the SpeechLM’s capability to cease speaking when interrupted.
@@
Downstream Evaluation. Downstream evaluation refers to evaluating the ability of SpeechLMs to perform specific tasks, such as ASR, TTS, Speaker Identification, etc. The evaluation can be performed on pre-trained models by adding few-shot example(s) at the start of the prompt or on the instruction-tuned models by directly instructing them to do so. Several benchmarks compile a broad array of downstream tasks to provide a comprehensive assessment of SpeechLMs. SUPERB [155] includes a variety of speech understanding tasks. SD-Eval [162] assesses the paralinguistic understanding of SpeechLMs using the emotion, age, environment, and age classification tasks. SALMON [165] tests the SpeechLMs’ ability to generate speech with consistent paralinguistic and environmental characteristics. Voicebench [166] evaluates SpeechLM general capabilities. Dynamic-SUPERB [163], MMAU [159], AirBench [161], and AudioBench [160] extend beyond traditional speech tasks to include sound and/or music-related challenges as well.
@@
While these benchmarks offer comprehensive coverage of audio-related tasks, they primarily require models to respond in text, which creates a barrier to end-to-end speech interaction evaluation. To address this limitation, VoxEval [167] focuses on benchmarking the knowledge understanding capabilities of SpeechLMs, providing question-answer pairs in comprehensive subjects in speech format, along with an evaluation pipeline tailored for speech output. Additionally, it presents questions under various input audio conditions for evaluating the model’s robustness and pioneers the spoken math reasoning evaluation.